# Articulation Work and Tinkering for Fairness in Machine Learning

### v1

@@ 4. Method and Data Analysis
Selbst et al. (2019) identify CS researchers working in fair AI as powerful. This is because the form a technology ultimately takes is shaped by the perspectives and practices of those that develop it: “as fair ML researchers seek to define the ‘best’ approach to fairness, we also implicitly decide which problems and relevant social groups are important to include in this process. Our choices prioritize certain views over others, exerting power in ways that must be accounted for” (Selbst et al., 2019). Not least for this reason, we study the gap between SOI and CS starting from the influential perspectives and practices of CS researchers.
@@
To our interview requests, we received eleven positive answers, with ten following through with the interview. The resulting group was therefore made up of established, fairly senior CS researchers, who are currently actively involved in fair AI research in Europe. Table 1 displays relevant information about our interview partners, such as the institutional position, country of institution, and gender.
@@
Figure 3. Sankey Diagram. Represented are 450 conference and workshop papers from 2009 to 2023 from our interview partners to give an overview of publication venues and research sub-area.
@@
The completed qualitative study follows an episodic narrative interview approach (Mueller, 2019). This method is used to improve the understanding of a phenomenon, in our case fair AI, by inviting the interview partners to tell their individual stories of life and work experience around said phenomenon. The interview format allows interview partners enough time for self-reflection and for the opportunity to share in richer detail their direct experiences. Collected data is thus useful to uncover the narratives around fair AI and related activities, which then require analysis and interpretation (Silverman, 2010). Our interviews were held throughout the year 2022, with the duration of each interview ranging from one to one and a half hours. Following every interview, the recordings were manually transcribed and pseudonymized. All personal identifiers were removed from our interview data and all our interview partners are referred to by pseudonyms i.e., R1, R2, etc. Subsequently, the interview material was interpreted using an explorative approach, informed by an inductive analysis that facilitates the possibilities of finding novel and unexpected issues (Corbin and Strauss, 1990). The first iterative phase of interpretation had been conducted closely to the material, resulting in hundreds of qualitative codes. Following this, we conducted the rest of the coding process, gradually clustering the codes into certain categories, merging, and rephrasing codes in numerous iterative phases. The resulting codes and sub-codes were reconsidered and collectively discussed many times.
@@ 5. Findings
Our findings are structured into three subsections. We start with two questions: First, what does the fair AI world want? Second, zoomed in on the CS lab, what do CS researchers do? Of course, the fair AI world also ‘does things’, and CS researchers also ‘want things’. Yet in this particular case, our focus lies on the (SOI) demands of the fair AI world to understand whether and how they align with what is done in the situated practice of the CS lab. In the last subsection, we further zoom into researchers’ micro-strategies for including and tinkering with SOI.
@@ 5.1. Social worlds: what does the fair AI world want?
In this section, we introduce the most prominent actors in the social world of fair AI that we identified in our interviews. We refer to it herein as the fair AI world. Our first finding is that there are three important collective actors: the CS community, project partners and regulatory bodies. These actors display an increasing orientation towards the SOI paradigm, but with different needs, expectations, and interpretations attached to it.
@@
The first significant actor that emerged in the interviews is the CS community. However, there was a time when our early-contributors to fair AI found it challenging to align their research with the then-research interests of the CS community. Particularly, the CS community was mainly concerned with research problems that tackled scalability or efficiency, as retold by R8, a computer science professor at an Italian university. It was, especially before 2015, that ‘fairness’ was neither a prominent research area in the CS community nor the widely popular topic it is today. R1, a senior scientist working at a German research lab, recalled that time as follows:
@@
Interests and demands of social worlds, such as in the case of the CS community, can shift and change (Knorr-Cetina, 1991, 1999). Now, that the CS community acknowledges research on fairness (metrics), the initial misalignment with the CS community “does not hurt anymore”, as to further quote R7. Early-contributors are now on good terms with their epistemic community, some of them also outlining how they played a crucial role in shaping the community’s evolving interests in fairness. Notably, the impact of their contributions is evident, if only taking into account citation figures. For instance, R9’s first fairness paper has over 1000 citations, while an early contribution by R8’s surpasses 4000 citations, at time of writing. There is a successful alignment of CS researchers’ work on fairness with the demands of the CS community, and their impact increases as the fair AI field grows.
@@
The second significant actor was the ‘project partner’. Typically, we found a project partner to be a private entity, and in our interviews, banks, credit agencies, and insurance companies were the ones most frequently mentioned. Public project partners were institutions such as hospitals, police, public broadcasters, or other research institutions. While there are many challenges for researchers in applied partnerships due to, for example, privacy concerns or licensing disputes (Holloway, 2015; Biscotti et al., 2012), private and public actors were perceived as particularly valuable because they can be a source of funding, interesting cases, datasets and models to study fairness. The latter was understood to be important for fair AI research because it required societal datasets containing information about people. A successful alignment with such data-providing partners was termed ‘real-world’ research, and demonstrating real-world impact was considered to be exceptionally relevant by our interview participants.
@@
“I actually see this as kind of paradigm shift. Several years back when I was a young researcher, ethical aspects in our kind of research were kind of an afterthought and a necessary evil but right now it’s kind of the first thing to consider and it’s always there” (R2).
@@
“In theory, we could have developed different models, one predicting the default within six months, one within one year and so on, so that every user could use the model for the period preferred. But this was cost-intense, e.g. building, maintaining several models, and so in the end it was not possible for the budget of the company to do that.” (R8)
@@
The majority of our interview partners anticipated advantages for their research once such regulations were implemented. For instance, R1 emphasized the necessity for “some common ground where you can stand on, and these are standards”. In a similar vein, R8 suggestd standardization bodies like the International Organization for Standardization (ISO) to control and audit AI development. He also proposed compensation mechanisms for “people who have been harmed by such systems”, and the introduction of general reporting mechanisms. These reporting systems would, as R8 explained, enable potentially disadvantaged individuals to denounce and record their cases.
Despite this predominantly expressed desire for standards and reporting mechanisms, at the time of the interviews, our participants mainly engaged with regulatory bodies through the perspective of their project partners. R8 noted how current regulatory attempts shifted his project partners’ interest towards identifying and mitigating algorithmic bias. He had observed a similar pattern in the past regarding (computational) privacy issues, wherein significant interest in the topic only emerged after the implementation of regulations.
“I have to say that the [company’s] interest in these [fairness] techniques was motivated by the rules, by the law. This is a general behaviour, for instance, [if you] think about privacy twenty years ago, nobody cared about privacy issues until the regulations came out.” (R8)
@@
What does the fair AI world want? We saw that the CS community’s research acknowledged fair AI research as important, project partners were perceived to shape and constrain research problems, and regulatory bodies were emerging with open consequences for social world actors, and CS researchers. As a result, we found most interview participants to be uncertain about how to interpret current dynamics and in which direction the fair AI world would ‘truly’ orient itself. As one interview participant, R1, slightly disillusioned, noted, “in the end, everyone has a different understanding of what is fair and what is not.” Against this backdrop, the expressed desire for regulatory standards can also be considered as an attempt to obtain guidance on how fairness can become a doable problem. Yet, simultaneous alignment with multiple demands of diverse social world actors increased articulation work. This is because their perceptions, requirements and demands were difficult to align.
@@ 5.2. Laboratories and experiments: what do CS researchers do?
At the level of the lab, our interview participants were often and understandably occupied with more pressing tasks and concerns that did not necessarily pertain to tending to the broader social claims that motivated their research. They needed to spend most of their time dividing their tasks within projects, acquiring new projects, and strategically pursuing new projects with partners to gain access to new resources. For instance, R5 conveyed that her current research projects on fair AI also offered secure employment opportunities.
@@
In a similar vein, fair AI projects were also a way to secure positions for preceding and younger researchers, and R6 remarked that the fair AI research projects he was currently working on were also important because they secured his students’ future positions in research. This is not to say that their research was not important to our interview participants, but, that on top of it, fair AI work had other dimensions beyond the undertaking of ethical and social justice-oriented claims. A fair AI laboratory is a workplace and also a site of career building for people. We witnessed that this reality manifested through practices such as placing staff on projects due to pre-existing working contracts or attempting to redefine work, to make it fit to existing knowledge and skills of the staff in the lab.
@@
Apart from shouldering the responsibility to keep the staff in their labs employed, some of our interviewed participants also saw how tending to that created an unexpected incongruity. Here, we refer to the call to integrate diversity and inclusion (D&I) considerations in AI systems, and how in pertaining to problems of unfairness and bias, these can be alleviated via diverse research teams (Shams et al., 2023). However, securing the mid to long-term positions of preexisting staff was seen to hinder these efforts. One interview participant was particularly concerned about this issue:
I think here we need to really think about how we can increase diversity, which is obviously important” (R4).
@@
Project-based funding mechanisms also provided a temporary constraint to the distribution of resources, as projects’ duration was always limited, typically spanning three to five years. Such time constraints therefore called for savvy project management skills, which would be better reflected towards the end of a project life cycle. Inadvertently, this led to prioritizing research problems that contributed to achieving project objectives on time, but that perhaps required “less effort” from an experimental setup and execution point of view. For instance, although some interview participants attempted to set up user studies in their fair AI research, they expressed dissatisfaction as the actual engagement with users remained limited due to time constraints. R2 emphasized the typicality of such situations.
@@
Yet, our interview participants often grappled with keeping up with the overwhelming amount of new scientific papers and information that emerged in the fair AI research area. When we asked about their primary tasks, we heard that a significant amount of time was allotted to following, reviewing, and reading the numerous scientific papers, to remain informed and up-to-date.
“Still one big point is really to figure out what other people do in this topic. It’s currently a hype topic and if it is a hype topic you have more publications than you can ever read.” (R2)
Articulation work encompassed not only the task of monitoring and reading the papers that were published but also the discernment of which papers were merely contributing to such ‘noise’, e.g. “go[ing] on and on about the same things” (R10), and those significant for fair AI research. In that regard, early-contributor R7 stated that “as soon as a particular few become hyped, over-hyped, many people jump on it and they aim to publish rather than to do something useful” (R7). The days of being an ‘exception’ for these CS researchers working on fairness research were long gone. In fact, due to the fast pace at which fair AI was evolving, CS researchers felt there were expectations to keep up with this high-speed publishing environment.
“I have to publish. I have to present research results. Of course, we have to be productive. But I think it’s good also to think a bit more, what we want to do in the end. You can have better publications, [but] I think we [should] have better contributions, actually.” (R5)
Interview participants, such as R5, showcased that producing publications was a requirement of sustaining or enhancing academic careers. Becoming part of the ‘noise’ was thus difficult to avoid, however, it also contributed to a lack of clarity about what constituted (or should constitute) impactful scientific progress for the fair AI world, even within the CS paradigm.
@@
We now focus on the production work of fair AI research, and related experiments. Early-contributors recounted that their first experiments in fair AI research were conducted on an ad-hoc basis, such as training a classifier on a biased dataset to experiment with the results. After obtaining results, they would convince their colleagues (and themselves) that fairness was a relevant research problem by putting a valuable object in the foreground: the (biased) dataset.
@@
Along with the dataset, other objects were placed in the foreground for production work. Over the past years, numerous fairness metrics and other bias mitigation techniques had been defined and developed, making it possible to “package” (Fujimura, 1987) them into fairness toolkits (Balayn et al., 2023). These collections of metrics, methods, and practical instructions were intended to ease their use in research, but more importantly, streamline their use in industry, where often inexperienced practitioners were the ones tasked with building and ‘de-biasing’ ML systems (Deng et al., 2022; Balayn et al., 2023). These objects thus facilitated alignment between the CS community and their project partners. And while the use of fairness metrics allowed for measuring, quantifying and comparing the output of models, at the same time, they could instill the idea that only what can be captured by such metrics should be defined and identified as fair (or unfair).
R9, a professor at a university in Belgium and another early-contributor to the field, emphasized that the primary objective of fair AI research should be to focus on model accuracy by “concentrat[ing] mainly on the distorted vision that is there [in the data used]”. Accuracy, for him, was part of addressing fundamental computational and statistical questions that the CS lab was supposed to pose, and not “political ones”. With this demarcation approach, interview participants like R9 suggested that CS can and should be more autonomous from socially-oriented research and political actors. While such interpretations of the role of technology could be considered clear deviations from an SOI approach, we recalled that a focus on datasets and on computationally relevant questions such as accuracy had allowed for alignment with the initially dismissive CS community. In this way, fair AI research could now be “stored” (Star, 1999) (e.g. in a dataset, in a computer, in a metric), shared across different levels of work organization, and made comparable. Such practices of storing, sharing and comparing, further facilitated alignment with other social world actors. For example, the creation of fairness metrics allowed for project partners to request that they be implemented into their own systems and provided a potential method to communicate standards adherence to regulatory bodies.
@@
What do CS researchers do? In this section, we witnessed that CS researchers were engaged in a substantive amount of articulation work: writing project proposals, securing staff positions, allocating time, reading all the papers and coming up with new ones. All this deviated researchers’ attention from engaging with fair AI research problems that they considered to be closer to SOI. As a result, the production work of fair AI research was mainly aligned with the CS community and CS positions.

### v2

@@ 4. Method and Data Analysis
Selbst et al. (2019) identify CS researchers working in fair AI as powerful. This is because the form a technology ultimately takes is shaped by the perspectives and practices of those that develop it: “as fair ML researchers seek to define the ‘best’ approach to fairness, we also implicitly decide which problems and relevant social groups are important to include in this process. Our choices prioritize certain views over others, exerting power in ways that must be accounted for” (Selbst et al., 2019). Not least for this reason, we study the gap between SOI and CS, starting from the influential perspectives and practices of CS researchers.
@@
To our interview requests, we received eleven positive answers, with ten following through with the interview. The resulting group was therefore made up of established, fairly senior, CS researchers, who are currently actively involved in fair AI research in Europe. Table 1 displays relevant information about our interview partners, such as the institutional position, country of institution, and gender.
@@
[…]
[A histogram was used to show the distribution of the h-index of the study.]Depicted is a histogram in blue colored boxes, with the y-axis showing the count, from 0 to 4 and the x-axis the value of the h-index, 5, 13, 21, 29, 37. There is one individual with an h-index between 5 and 13, 2 individuals with an h-index ranging from 13 and 21, 4 individuals with a value between 21 and 29, 2 individuals with an h-index value between 20 and 37, and one individual with an h-index value higher than 37.
@@
[…]
[A sankey diagram was used to give an overview of the publication venues and research sub-areas where our interview partners participate (last five years).]A sankey diagram illustrates the relationships between publishing venue and research sub-area based on 450 research papers belonging to the 10 interview partners. On the left side of the diagram, there are 11 names for conferences and workshops. From top to bottom: CEUR workshop proceedings, ACM international conference proceeding series, proceedings of the ACM SIGKDD conference, IEEEE international conference on data mining, International conference on World Wide Web, proceedings of machine learning research, eye tracking research and applications symposium, content-based multimedia indexing, issues on sentiment discovery and opinion mining, ACM symposium on applied computing, conference on human factors in computing systems. This site is then linked to the middle column containing the research sub-area by colored, curved lines. From the top: social networking, artificial intelligence, learning systems, data handling, ontology, hypertext systems, multimedia contents, hypermedia systems, education, data description, semantics, adaptive systems, visual content, semantic web, data mining, world wide web, performance, machine learning, eye tracking, eye movements, indexing of information, sentiment analysis, computation theory, human engineering. These items are all joined with the column on the right, which represents the language of the published works (English). They are also joined with curved, colored lines that differ in thickness depending on the number of research papers.
Figure 3. Sankey Diagram. Represented are 450 conference and workshop papers from 2009 to 2023 from our interview partners to give an overview of publication venues and research sub-areas.
@@
The completed qualitative study follows an episodic narrative interview approach (Mueller, 2019). This method is used to improve the understanding of a phenomenon, in our case fair AI, by inviting the interview partners to tell their individual stories of life and work experience around said phenomenon. The interview format allows interview partners enough time for self-reflection and for the opportunity to share in richer detail their direct experiences. The collected data is thus useful to uncover the narratives around fair AI and related activities, which then require analysis and interpretation (Silverman, 2010). Our interviews were held throughout the year 2022, with the duration of each interview ranging from one to one and a half hours. Following every interview, the recordings were manually transcribed and pseudonymized. All personal identifiers were removed from our interview data and all our interview partners are referred to by pseudonyms i.e., R1, R2, etc. Subsequently, the interview material was interpreted using an explorative approach, informed by an inductive analysis that facilitates the possibilities of finding novel and unexpected issues (Corbin and Strauss, 1990). The first iterative phase of interpretation had been conducted closely to the material, resulting in hundreds of qualitative codes. Following this, we conducted the rest of the coding process, gradually clustering the codes into certain categories, merging, and rephrasing codes in numerous iterative phases. The resulting codes and sub-codes were reconsidered and collectively discussed many times.
@@ 5. Findings
Our findings are structured into three subsections. We start with two questions: First, what does the fair AI world want? Second, zoomed in on the CS lab, what do CS researchers do? Of course, the fair AI world also ‘does things’, and CS researchers also ‘want things’. Yet in this particular case, our focus lies on the (SOI) demands of the fair AI world to understand whether and how they align with what is done in the situated practice of the CS lab. In the last subsection, we further zoom into researchers’ micro-strategies for including and tinkering with SOI demands.
@@ 5.1. Social worlds: what does the fair AI world want?
In this section, we introduce the most prominent actors in the social world of fair AI that we identified in our interviews. We refer to it herein as the fair AI world. Our first finding is that there are three important collective actors: the CS community, project partners, and regulatory bodies. These actors display an increasing orientation towards the SOI paradigm but with different needs, expectations, and interpretations attached to it.
@@
The first significant actor that emerged in the interviews was the CS community. However, there was a time when our early contributors to fair AI found it challenging to align their research with the then-research interests of the CS community. Particularly, the CS community was mainly concerned with research problems that tackled scalability or efficiency, as retold by R8, a computer science professor at an Italian university. It was, especially before 2015, that ‘fairness’ was neither a prominent research area in the CS community nor the widely popular topic it is today. R1, a senior scientist working at a German research lab, recalled that time as follows:
@@
Interests and demands of social worlds, such as in the case of the CS community, can shift and change (Knorr-Cetina, 1991, 1999). Once the CS community acknowledged research on fairness (metrics), the initial misalignment with the CS community “does not hurt anymore”, to further quote R7. Early contributors also outlined how they played a crucial role in shaping the community’s evolving interests in fairness. Notably, the impact of their contributions is evident, if only taking into account citation figures. For instance, R9’s first fairness paper has over 1000 citations, while an early contribution by R8 surpasses 4000 citations, at the time of writing. There is a successful alignment of CS researchers’ work on fairness with the demands of the CS community, and their impact increases as the fair AI field grows.
@@
The second significant actor was the ‘project partner’. Typically, we found a project partner to be a private entity, and in our interviews, banks, credit agencies, and insurance companies were the ones most frequently mentioned. Public project partners were institutions such as hospitals, police, public broadcasters, or other research institutions. While there are many challenges for researchers in applied partnerships due to, for example, privacy concerns or licensing disputes (Holloway, 2015; Biscotti et al., 2012), private and public actors were perceived as particularly valuable because they can be a source of funding, interesting use cases, and datasets and models to study fairness. The latter was understood to be important for fair AI research because it required societal datasets containing information about people. A successful alignment with such data-providing partners was termed ‘real-world’ research, and demonstrating real-world impact was considered to be exceptionally relevant by our interview participants.
@@
“I actually see this as a kind of paradigm shift. Several years back when I was a young researcher, ethical aspects in our kind of research were kind of an afterthought and a necessary evil but right now it’s kind of the first thing to consider and it’s always there” (R2).
@@
“In theory, we could have developed different models, one predicting the default within six months, one within one year, and so on, so that every user could use the model for the period preferred. But this was cost-intense, e.g. building, maintaining several models, and so in the end it was not possible for the budget of the company to do that.” (R8)
@@
The majority of our interview partners anticipated advantages for their research once such regulations were implemented. For instance, R1 emphasized the necessity for “some common ground where you can stand on, and these are standards”. In a similar vein, R8 suggested standardization bodies like the International Organization for Standardization (ISO) control and audit AI development. He also proposed compensation mechanisms for “people who have been harmed by such systems”, and the introduction of general reporting mechanisms. These reporting systems would, as R8 explained, enable potentially disadvantaged individuals to denounce and record their cases.
Despite this predominantly expressed desire for standards and reporting mechanisms, at the time of the interviews, our participants mainly engaged with regulatory bodies through the perspective of their project partners. R8 noted how current regulatory attempts shifted his project partners’ interest toward identifying and mitigating algorithmic bias. He had observed a similar pattern in the past regarding (computational) privacy issues, wherein significant interest in the topic only emerged after the implementation of regulations.
“I have to say that the [company’s] interest in these [fairness] techniques was motivated by the rules, by the law. This is a general behavior, for instance, [if you] think about privacy twenty years ago, nobody cared about privacy issues until the regulations came out.” (R8)
@@
What does the fair AI world want? We saw that the CS community’s research acknowledged fair AI research as important, project partners were perceived to shape and constrain research problems, and regulatory bodies were emerging with open consequences for social world actors, and CS researchers. As a result, we found most interview participants to be uncertain about how to interpret current dynamics and in which direction the fair AI world would ‘truly’ orient itself. As one interview participant, R1, slightly disillusioned, noted, “in the end, everyone has a different understanding of what is fair and what is not.” Against this backdrop, the expressed desire for regulatory standards can also be considered as an attempt to obtain guidance on how fairness can become a doable problem. Yet, simultaneous alignment with multiple demands of diverse social world actors increased articulation work. This is because their perceptions, requirements, and demands were difficult to align.
@@ 5.2. Laboratories and experiments: what do CS researchers do?
At the level of the lab, our interview participants were often and understandably occupied with more pressing tasks and concerns that did not necessarily pertain to tending to the broader social claims that motivated their research. They needed to spend most of their time dividing their tasks within and between projects, acquiring new projects, and strategically pursuing new projects with partners to gain access to new resources. For instance, R5 conveyed that her current research projects on fair AI also offered secure employment opportunities.
@@
In a similar vein, fair AI projects were also a way to secure positions for preceding and younger researchers, as R6 remarked that the fair AI research projects he was currently working on were also important because they secured his students’ future positions in research. This is not to say that their research was not important to our interview participants, but, that on top of it, fair AI work had other dimensions beyond the undertaking of ethical and social justice-oriented claims. A fair AI laboratory is a workplace and also a site of career building for people. We witnessed that this reality manifested through practices such as placing staff on projects due to pre-existing working contracts or attempting to redefine work, to make it fit to existing knowledge and skills of the staff in the lab.
@@
Some of our interviewed participants also saw how shouldering the responsibility to keep the staff in their labs employed created an unexpected incongruity. Here, we refer to the call to integrate diversity and inclusion (D&I) considerations in AI systems and the claim that problems of unfairness and bias can be alleviated via diverse research teams (Shams et al., 2023). However, securing the mid to long-term positions of preexisting staff was seen to hinder these efforts. One interview participant was particularly concerned about this issue:
“I think here we need to really think about how we can increase diversity, which is obviously important” (R4).
@@
Project-based funding mechanisms also provided a temporary constraint to the distribution of resources, as project duration was always limited, typically spanning three to five years. Such time constraints therefore called for savvy project management skills, particularly towards the end of a project life cycle. Inadvertently, this led to prioritizing research problems that contributed to achieving project objectives on time, but that perhaps required “less effort” from an experimental setup and execution point of view. For instance, although some interview participants attempted to set up user studies in their fair AI research, they expressed dissatisfaction as the actual engagement with users remained limited due to time constraints. R2 emphasized the typicality of such situations.
@@
Yet, our interview participants often grappled with keeping up with the overwhelming amount of new scientific papers and information that emerged in the fair AI research area. When we asked about their primary tasks, we heard that a significant amount of time was allotted to following, reviewing, and reading the numerous scientific papers being published, to remain informed and up-to-date.
“Still, one big point is really to figure out what other people do on this topic. It’s currently a hype topic and if it is a hype topic you have more publications than you can ever read.” (R2)
Articulation work encompassed not only the task of monitoring and reading new papers, but also the discernment of which papers were merely contributing to such ‘noise’, e.g. “go[ing] on and on about the same things” (R10), and those significant for fair AI research. In that regard, early-contributor R7 stated that “as soon as a particular few become hyped, over-hyped, many people jump on it and they aim to publish rather than to do something useful” (R7). The days of these researchers being an ‘exception’ for working on fairness research were long gone. In fact, due to the fast pace at which fair AI was evolving, CS researchers felt there were expectations to keep up with this high-speed publishing environment.
“I have to publish. I have to present the research results. Of course, we have to be productive. But I think it’s good also to think a bit more, about what we want to do in the end. You can have better publications, [but] I think we [should] have better contributions, actually.” (R5)
Interview participants, such as R5, showcased that producing publications was a requirement for sustaining or enhancing academic careers. Becoming part of the ‘noise’ was thus difficult to avoid, however, it also contributed to a lack of clarity about what constituted (or should constitute) impactful scientific progress for the fair AI world, even within the CS paradigm.
@@
We now focus on the production work of fair AI research and related experiments. Early contributors recounted that their first experiments in fair AI research were conducted on an ad-hoc basis, such as training a classifier on a biased dataset to experiment with the results. After obtaining results, they would convince their colleagues (and themselves) that fairness was a relevant research problem by putting a valuable object in the foreground: the (biased) dataset.
@@
Along with the dataset, other objects were placed in the foreground for production work. Over the past years, numerous fairness metrics and other bias mitigation techniques have been defined and developed, making it possible to “package” (Fujimura, 1987) them into fairness toolkits (Balayn et al., 2023). These collections of metrics, methods, and practical instructions were intended to ease their use in research, but more importantly, streamline their use in industry, where often inexperienced practitioners were the ones tasked with building and ‘de-biasing’ ML systems (Deng et al., 2022; Balayn et al., 2023). These objects thus facilitated alignment between the CS community and their project partners. And while the use of fairness metrics allowed for measuring, quantifying, and comparing the output of models, at the same time, they could instill the idea that only what can be captured by such metrics should be defined and identified as fair (or unfair).
R9, a professor at a university in Belgium and another early contributor to the field, emphasized that the primary objective of fair AI research should be to focus on model accuracy by “concentrat[ing] mainly on the distorted vision that is there [in the data used]”. Accuracy, for him, was part of addressing fundamental computational and statistical questions that the CS lab was supposed to pose, and not “political ones”. With this demarcation approach, interview participants like R9 suggested that CS can and should be more autonomous from socially-oriented research and political actors. While such interpretations of the role of technology could be considered clear deviations from an SOI approach, we recalled that a focus on datasets and on computationally relevant questions such as accuracy had allowed for alignment with the initially dismissive CS community. In this way, fair AI research could now be “stored” (Star, 1999) (e.g. in a dataset, in a computer, in a metric), shared across different levels of work organization, and made comparable. Such practices of storing, sharing, and comparing, further facilitated alignment with other social world actors. For example, the creation of fairness metrics allowed for project partners to request that they be implemented into their own systems and provided a potential method to communicate standards adherence to regulatory bodies.
@@
What do CS researchers do? In this section, we witnessed that CS researchers were engaged in a substantive amount of articulation work: writing project proposals, securing staff positions, allocating time, reading emerging papers, and coming up with new ones. All this deviated researchers’ attention from engaging with fair AI research problems that they considered to be closer to the SOI paradigm. As a result, the production work of fair AI research was mainly aligned with the CS community and CS positions.